OpenStudy (anonymous):

if y=4 when x=-2, find x when y=6

OpenStudy (anonymous):

-2/4 =x/6 -2x=24 x=?

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

24/-2=-12

OpenStudy (anonymous):

dont you multiply criss cross ways

OpenStudy (anonymous):

Wait... you're right because i set it up wrong

OpenStudy (anonymous):

Mmhmm. You were right. It was supposed to be -12=4x

OpenStudy (anonymous):

y varies directly as xsquared : If y =15 when x=5 , find y when x =10

OpenStudy (anonymous):

are u still there

OpenStudy (anonymous):

y varies directly as xsquared : If y =15 when x=5 , find y when x =10

OpenStudy (anonymous):

im here sorry i's hectic over here

OpenStudy (anonymous):

yx^2=k, right? 15(25)=k 375=k So then 375=y10^2 375=100y y=3.75

OpenStudy (anonymous):

Double check?

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

\When two variables vary directly, it means they're multiplied and equal k (constant)

OpenStudy (anonymous):

So y times x^2 equals k We plug in the numbers (15)(5^2)=k (15)(25)=k k=375

OpenStudy (anonymous):

Good so far?

OpenStudy (anonymous):

what about the find y when x =10

OpenStudy (anonymous):

y varies inversely as x: If x=8 when y=9 , find x when y=18

OpenStudy (anonymous):

When y equals ten, you plug y and k back into the original equation, so 10^2y=375 y=3.75

OpenStudy (anonymous):

WAIT NO TAT's BACKWARDS

OpenStudy (anonymous):

15/25=k .6=k

OpenStudy (anonymous):

100y=.6 y=.006

OpenStudy (anonymous):

So inversely means multiplied. 9(8)=k k=72 18x=72 x=4

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Okay, where did I lose you?

OpenStudy (anonymous):

ithought that i would multiply the 8 x 18 and the 9 and x together

OpenStudy (anonymous):

Any variation equals a constant. So xy=k or x/y=k or x/yz=k

OpenStudy (anonymous):

Directly means divide, so x/y=k What were the variables again?

OpenStudy (anonymous):

x=8 when y =9, find x when y=18

OpenStudy (anonymous):

Okay, an this is inversely, right? So xy=k You plug in x and y and get (8)(9)=k 72=k

OpenStudy (anonymous):

Now that you have the constant, you plug it back into the original equation. xy=72

OpenStudy (anonymous):

You also have y now, so plug that in too for 18x=72 Now solve. x=72/18 x=4

OpenStudy (anonymous):

so y=4

OpenStudy (anonymous):

Close. y was 18, so x is 4

OpenStudy (anonymous):

if x=2 when y=10, find y when x=4

OpenStudy (anonymous):

how do they vary?

OpenStudy (anonymous):

oh, i got it. sorry. Okay, so 2/10=4/y 40=2y y=?

OpenStudy (anonymous):

xy=k 2(10)=k k=20 4y=20 y=5

OpenStudy (anonymous):

Yeah you were right. Sorry, i tried to set it up in fractions.